A company is considering a three-year project. New equipment will cost $200,000. The equipment falls in the MACRS three-year class (.3333, .4445, .1481, .0741) it will have a salvage value at the end of the project of $50,000. The project is expected to produce sales of $100,000 in the first year and sales will increase by $50,000 each year after that. Expense are expected to be 40% of sales. An investment in net-working capital of $5,000 is required at the beginning of the project. In year 2, net working captial must be increased by $10,000. All net-working capital will be recovered in the final year of the project. The tax rate is 30%. The company is funded with 40% debt with a yield of 7% and 60% equity with a required return of 15%.

A. What are the annual depreciation expenses and the after-tax salvage value of the machine? (ATSV)

B. What is the Operating Cash Flow for each of the three years?

C. What are the total cash flows for years 0-3?

D. What is the company's hurdle rate?

E. What is the NPV, IRR, and Profitability Index? Show work please.

F. Should the company invest in this project and why?
Answer #1

a). Annual depreciation expenses are:

Formula Year (n) 1 2 3 4
Asset cost 200000 200000 200000 200000
%age depreciation 33.33% 44.45% 14.81% 7.41%
Asset cost*%age dep. Depreciation expense (D) 66660 88900 29620 14820

After-tax salvage value = salvage value - tax rate*(salvage value - book value)

= 50,000 - 30%*(50,000 - 14,820) = 39,446

d). Hurdle rate for the company (or WACC) = (weight of debt*cost of debt*(1-Tax rate)) + (weight of equity*cost of equity)

= (40%*7%*(1-30%)) + (60%*15%) = 10.96%

Answers Formula Year (n) 0 1 2 3
Initial investment (II) 200000
Sn-1 + 50,000 Sales (S) 100000 150000 200000
40%*S Expenses (E) 40000 60000 80000
Depreciation (D) 66660 88900 29620
S-E-D EBIT -6660 1100 90380
30%*EBIT Tax @ 30% -1998 330 27114
EBIT-Tax Net income (NI) -4662 770 63266
Add: Depreciation (D) 66660 88900 29620
Part (b) NI+D Operating Cash Flow (OCF) 61998 89670 92886
Less: Inc. in NWC -5000 0 -10000 15000
Add: After-tax salvage value (ASV) 39446
Part (c) OCF - Inc. in NWC + ASV - II Free Cash Flow (FCF) -205000 61998 79670 147332
1/(1+10.96%)^n Discount factor @ 10.96% 1.000 0.901 0.812 0.732
FCF*Discount factor PV of FCF -205000.00 55874.19 64708.59 107844.44
Part (e) Sum of all PVs NPV 23427.21
Part (e) Using IRR function with FCFs IRR 16.53%
Part (e) Sum of PV of all cash inflows/initial cash outflow Profitability index (PI) 1.11

f). The company should invest in this project because it has a positive NPV and its IRR is greater than the company's hurdle rate.